# Maximum Implementable Improvements over Deferred Acceptance

### v2

@@ 4.1 Larger Improvements with Less Justified Envy
We find Corollary 1 interesting because DA+TTC has been considered suboptimal regarding a number of axioms (Kesten, 2010; Troyan et al., 2020), yet it is not uniformly doubly dominated by the more well-regarded EADA mechanism.
4.2 Consent and Improvement Size
In his seminal paper, Kesten (2010) argues that “the likelihood of getting stuck without gaining much Pareto improvement clearly increases with the cycle size". This intuition seems natural—coordinating more students should be harder. However, this relationship is more nuanced than originally thought.
Let us recall that, in EADA, a student agrees to waive his priorities before the DA mechanism is executed at schools where he is an interrupter. Kesten shows that such students never benefit nor gain from giving their consent. Let us divide students who consent into two classes: non-beneficiary consenters are those who neither benefit nor are harmed by giving their consent, and beneficiary consenters who explicitly benefit from the trading cycles generated by their consent to waive their priorities. It is reasonable to conjecture that beneficiary consents are more likely to be obtained than non-beneficiary ones, as the former ones directly benefit rather than providing consent altruistically.
With this distinction in mind, let us revisit Example 2 in the proof of Proposition 1. EADA implements the four-cycle ($i_{1}\to i_{6}\to i_{4}\to i_{5}\to i_{1}$), requiring non-beneficiary consent from both $i_{3}$ (at $s_{6}$) and $i_{7}$ (at $s_{4}$), as well as beneficiary consent from $i_{5}$ (at $s_{6}$). In contrast, the maximum improvement implements two disjoint cycles requiring fewer non-beneficiary consents (1 versus 2 of EADA), and the same number of beneficiary consents (one each, see Table 1). If non-beneficiary consents, which rely purely on altruism, fail, EADA cannot improve any student, and yet the largest improvement can still execute the two-cycle ($i_{1},i_{2},i_{1}$), given the consent of potential beneficiary $i_{5}$. Student $i_{5}$ does not benefit given the lack of non-beneficiary consents, but since consents are expressed ex-ante this does not affect his decision on whether to consent or not.
Table 1: Consent Requirements in Example 2
Mechanism Consent Required Improvement Size
Non-beneficiary Beneficiary
EADA $i_{3},i_{7}$ $i_{5}$ 4 ($i_{1}\to i_{6}\to i_{4}\to i_{5}\to i_{1}$)
$i_{4}$, $i_{3}$ (no $i_{7}$) 2 ($i_{1}\to i_{5}\to i_{1}$)
$i_{5}$ (no $i_{7},i_{4}$) 2 ($i_{1}\to i_{2}\to i_{1}$)
otherwise 0
Maximum $i_{7}$ $i_{1}$ 6 ($i_{1}\to i_{2}\to i_{1}$), ($i_{3}\to i_{6}\to i_{4}\to i_{5}\to i_{3}$)
Improvement $i_{5}$ 2 ($i_{1}\to i_{2}\to i_{1}$)
otherwise 0
We summarize this observation as Proposition 2 below.
Proposition 2.
There exist problems for which a Pareto improvement that strictly improves more students than EADA requires fewer total and non-beneficiary consents.
Furthermore, in some problems, larger improvements can be implemented relying solely on beneficiary consents, whereas EADA may fail to implement smaller improvements without non-beneficiary consent.
Having examined the potential shortcomings of existing mechanisms, we proceed to analyse a new mechanism that improves as many students as possible compared to the DA baseline.
5 Maximum Implementable Improvement over Deferred Acceptance
Our theoretical analysis has revealed that existing mechanisms can fall short of maximizing the number of students who benefit from efficiency enhancements to DA. The Maximum Implementable Improvement over Deferred Acceptance (MIIDA) mechanism addresses this limitation by identifying and implementing improvements that benefit the maximum possible number of students given the consent structure.
To define MIIDA formally, let us incorporate consent into the mathematical framework. Each student reports strict preferences over schools $\succ_{i}$ and whether they consent $c_{i}=1$ (or not, with $c_{i}=0$) to waive their priority at schools. A consent-based mechanism maps a school choice problem and a consent decision by each student to a matching. Let $C\subseteq I$ denote the set of students who agree to waive their priority when they do not benefit from the proposed trades. For any edge $i\to j$ in the envy digraph $G^{\mathtt{DA}(P)}$, let $\text{consent}(i\to j)$ denote the set of students whose consent is required for student $i$ to receive student $j$’s DA assignment. Specifically,
$\text{consent}(i\to j)=\{k\in I:k\text{ has higher priority than }i\text{ at school }\mathtt{DA}_{j}(P)\}$
For any feedback set $F$, we define $L(F)=\bigcup_{C\in F}\bigcup_{(i\to j)\in C}\text{consent}(i\to j)$ as the set of all students whose consent is required to implement the trades in $F$. A feedback set $F$ is implementable given consent set $C$ if $L(F)\setminus V(F)\subseteq C$, so that all required consents come from either beneficiaries of the trades or students in the consent set $C$.
We define MIIDA as follows.
Algorithm 1 Maximum Improvement over Deferred Acceptance (MIIDA)
1:Compute the student-optimal stable matching $\mu^{0}=\mathtt{DA}(P)$
2:Set iteration counter $t=0$
3:repeat
4:     Generate the envy digraph $G^{\mu^{t}}$ for current matching $\mu^{t}$
5:     if $G^{\mu^{t}}$ contains no cycles then
6:         break
7:     end if
8:     Let $N^{t}=I\setminus C$ be the set of non-consenting students
9:     Remove nodes $N^{t}$ and adjacent edges from $G^{\mu^{t}}$ to obtain reduced graph $\tilde{G}^{\mu^{t}}$
10:     if $\tilde{G}^{\mu^{t}}$ contains no cycles then
11:         break $\triangleright$ No improvements possible with available consent
12:     end if
13:     Find maximum feedback set $F^{t}$ in $\tilde{G}^{\mu^{t}}$ (with a fixed tie-breaking rule)
14:     Implement trades in $F^{t}$ to produce matching $\mu^{t+1}$
15:     $t\leftarrow t+1$
16:until convergence
17:return final matching $\mu^{t}$
MIIDA iteratively finds maximum implementable feedback sets in the envy digraph until no further improvements are possible. Iteration is essential because implementing trades changes the envy structure: after students exchange assignments, new envy relationships emerge while others disappear. To illustrate, consider the consent-labeled envy digraph from our earlier example. MIIDA proceeds as follows:
  1. 1.
    In the first iteration, MIIDA identifies the maximum feedback set
    $F={(i_{1}\to i_{2}\to i_{1}),(i_{3}\to i_{6}\to i_{4}\to i_{5}\to i_{3})}$
    covering 6 students, which requires consent from $L(F)\setminus V(F)={i_{7}}$.
  2. 2.
    If $\{i_{1},i_{7}\}\subset C$, MIIDA implements $F$.
  3. 3.
    If the required consents are not available, MIIDA seeks the largest implementable feedback set in the graph. If $i_{7}\in C$, it finds $F^{\prime}={(i_{1}\to i_{2}\to i_{1})}$ covering 2 students.
  4. 4.
    If $i_{5}\in C$, implement $F^{\prime}$. Otherwise, make no changes over DA.
MIIDA satisfies a number of important desiderata.
Theorem 2.
MIIDA has the following properties:
  1. 1.
    Weakly Pareto-dominates DA.
  2. 2.
    Is Pareto-efficient if all students consent.
  3. 3.
    Only violates the priorities of students who consent.
  4. 4.
    Maximises the number of students who benefit from Pareto improvements over DA given the consent structure (so it is never doubly-dominated).
  5. 5.
    Can generate strictly fewer, but also strictly more, blocking pairs than EADA and DA+TTC, depending on the school choice problem.
  6. 6.
    No student has incentives not to consent. Specifically, for every student consenting is a very weakly dominant strategy in the complete information preference revelation game induced by MIIDA in which students simultaneously report their preferences and whether they consent to waive their priorities. [^7] A strategy is very weakly dominant for a player if, for every strategy profile, it gives a weakly higher utility than every other strategy (e.g. Chen and Micali, 2015). On the other hand, a weakly dominant strategy additionally requires that for some strategy profile it must yield a strictly higher utility than any other strategy available to the player.
  7. 7.
    Can be computed in polynomial time.
Proof.
  1. 1.
    Pareto-dominance: Follows directly from MIIDA’s construction: it implements only cycles where each participant receives a strictly preferred school relative to DA (by envy digraph definition). Non-participants keep their DA assignments, guaranteeing weak Pareto dominance for everyone.
  2. 2.
    Pareto-efficient if all consent: When $C=I$, $\tilde{G}^{\mu^{t}}=G^{\mu^{t}}$. Termination occurs only when no cycles exist, implying no further Pareto improvements are possible (by Lemma 1)).
  3. 3.
    Priority violations require consent: For any edge $i\to j$ implemented, consent is required from all $k$ with higher priority at $s=\mathtt{DA}_{j}(P)$. The condition $L(F^{t})\setminus V(F^{t})\subseteq C$ ensures only consenting non-beneficiaries’ priorities are violated.
  4. 4.
    Maximizes improvements given consent: Let $S$ be students improved by MIIDA. Suppose another mechanism $M$ improves $T$ with $|T|>|S|$. By Lemma 1, $T$ corresponds to disjoint cycles $\mathcal{C}_{T}$ in $G^{\mathtt{DA}(P)}$ implementable under $C$. In MIIDA’s first iteration, it selects a feedback set $F^{0}$ maximizing $|V(F^{0})|$ over implementable sets. Since $\mathcal{C}_{T}$ is implementable: $|V(F^{0})|\geq|T|$. But $S$ includes $V(F^{0})$ (at minimum), so $|S|\geq|T|$, contradiction. Thus no such $M$ exists.
  5. 5.
    Variable blocking pairs: Proven by Theorem 1 and Proposition 1: Example 1 shows MIIDA can generate more blocking pairs than EADA/DA+TTC; Example 2 shows it can also generate fewer.
  6. 6.
    Consent incentives: Fix student $i\in I$, any preference profile $P=(\succ^{\prime}_{i},\succ^{\prime}_{-i})$, and any consent decisions $c_{-i}$ by other students. Let $\mu^{\mathtt{DA}(P)}=\mathtt{DA}(P)$ and let $C=\{j\neq i:c_{j}=1\}\cup\{i\}$ denote the consent set when student $i$ consents. Let $F$ denote the maximum implementable feedback set under consent set $C$.
    When $c_{i}=0$, by MIIDA’s construction, $i$ cannot participate in any trading cycle. Therefore $\mathtt{MIIDA}_{i}(P;c_{i}=0,c_{-i})=\mu^{\mathtt{DA}(P)}_{i}$.
    When $c_{i}=1$:
    • If $i\notin V(F)=\mathtt{MIIDA}_{i}(P;c_{i}=1,c_{-i})=\mu^{\mathtt{DA}(P)}_{i}$
    • If $i\in V(F)=\mathtt{MIIDA}_{i}(P;c_{i}=1,c_{-i})=\mu^{\mathtt{DA}(P)}_{\text{succ}(i)}$
    where $\text{succ}(i)$ is $i$’s successor in the cycle containing $i$ within $F$. If $i\in V(F)$, then by definition of envy digraph $G^{\mu^{\mathtt{DA}(P)}}$, there exists edge $i\to\text{succ}(i)$, which means $\mu^{\mathtt{DA}(P)}_{\text{succ}(i)}\succ^{\prime}_{i}\mu^{\mathtt{DA}(P)}_{i}$. Therefore $\textup{rk}_{i}(\mu^{\mathtt{DA}(P)}_{\text{succ}(i)})<\textup{rk}_{i}(\mu^{\mathtt{DA}(P)}_{i})$.
    It follows that
    $\textup{rk}_{i}(\mathtt{MIIDA}(P;c_{i}=1,c_{-i}))\leq\textup{rk}_{i}(\mu^{\mathtt{DA}(P)}_{i})=\textup{rk}_{i}(\mathtt{MIIDA}(P;c_{i}=0,c_{-i}))$
    with strict inequality when $i\in V(F)$. Since this holds for arbitrary $(\succ^{\prime}_{i},\succ^{\prime}_{-i},c_{-i})$, consenting is a very weakly dominant strategy for student $i$.
  7. 7.
    Polynomial runtime: MIIDA runs in polynomial time through its iterative structure. DA is computed once initially in $O(n^{2})$ time (Gale and Shapley, 1962). Each subsequent iteration involves: (i) generating the envy digraph for the current matching in $O(n^{2})$ time, and (ii) finding a maximum feedback set. The cycle packing problem can be solved in $O(m\sqrt{n})$ time where $m$ is the number of edges and $n$ is the number of vertices (Biro et al., 2009). Since the envy digraph has at most $n^{2}$ edges, this step takes $O(n^{2}\sqrt{n})=O(n^{2.5})$ time per iteration. Since each iteration covers at least one additional student in trading cycles, and at most $n-1$ students can be improved over DA (Tang and Yu, 2014), MIIDA terminates within $n$ iterations. Therefore, the overall complexity is $O(n^{3.5})$.
While MIIDA improves as many students as possible given the consent structure available, and is never doubly dominated, it also has a number of important limitations. Arguably, the most severe is its lack of stability. Since it chooses cycles without taking into account priorities, it can lead to more blocking pairs than EADA and DA+TTC (but also fewer, as we have seen). In fact, MIIDA sometimes generates a set of blocking pairs that is a strict superset of EADA’s blocking pairs. Unfortunately, this is unavoidable by any mechanism that does not miss a significant number of improvement cycles.
To state this result formally, let us denote by $B(M,P)$ the set of blocking pairs generated by mechanism $M$ in problem $P$. We say mechanism $M$ is setwise justified envy minimal if for any problem $P$, there is no alternative mechanism $M^{\prime}$ such that $B(M^{\prime},P)\subsetneq B(M,P)$. A number of mechanisms, including EADA (Doğan and Ehlers, 2021a), satisfy this partial comparison property. However, setwise justified envy minimality is at odds with larger improvements.
Theorem 3.
Among Pareto-efficient mechanisms that weakly Pareto-dominate DA, every mechanism that is setwise justified envy minimal achieves an improvement ratio of $\frac{n-1}{2}$.
Proof.
Recall Example 1 in the proof of Theorem 1. In the case with $7$ students and schools, there are six cycles in DA’s envy digraph, summarized in Table 2.
Table 2: Cycles and Blocking Pairs in Example 1
Cycles Blocking Pairs
$K_{1}=(i_{1}\to i_{2}\to i_{1})$ $(i_{7},s_{1})$
$K_{2}=(i_{1}\to i_{2}\to i_{3}\to i_{1})$ $(i_{7},s_{1}),(i_{6},s_{1}),(i_{2},s_{1})$
$K_{3}=(i_{1}\to i_{2}\to i_{3}\to i_{4}\to i_{1})$ $(i_{7},s_{1}),(i_{6},s_{1}),(i_{2},s_{1})$
$K_{4}=(i_{1}\to i_{2}\to i_{3}\to i_{4}\to i_{5}\to i_{1})$ $(i_{7},s_{1}),(i_{6},s_{1}),(i_{2},s_{1})$
$K_{5}=(i_{2}\to i_{3}\to i_{4}\to i_{5}\to i_{2})$ $(i_{1},s_{2})$
$K_{6}=(i_{1}\to i_{6}\to i_{1})$ $(i_{7},s_{1}),(i_{2},s_{1}),(i_{5},s_{6})$
$K_{7}=(i_{1}\to i_{6}\to i_{2}\to i_{1})$ $(i_{7},s_{1}),(i_{1},s_{2}),(i_{5},s_{6})$
Any efficient mechanism that weakly dominates DA must implement at least one of these cycles (potentially two if mutually disjoint). In particular, MIIDA implements ($i_{1}\to i_{6}\to i_{1}$) and ($i_{2}\to i_{3}\to i_{4}\to i_{5}\to i_{2}$).
The cycles $K_{2},K_{3},K_{4},K_{6},K_{7}$ all generate a set of blocking pairs that is strictly larger (by set inclusion) than that generated by $K_{1}$. The cycle $K_{5}$ is incomparable to $K_{1}$, but implementing it by itself does not lead to a Pareto-efficient allocation, since the cycle $K_{6}$ could be jointly implemented. But if both $K_{5}$ and $K_{6}$ are implemented, the set of blocking pairs becomes $\{(i_{1},s_{2}),(i_{2},s_{1}),(i_{7},s_{1})\}$, a strict superset of the set of blocking pairs generated by $K_{1}$. Thus, we conclude that every setwise justified envy minimal mechanism must implement $K_{1}$therefore missing the larger improvements dictated by $K_{5}$ and $K_{6}$ jointly, leading to an improvement ratio of $\frac{n-1}{2}=\frac{6}{2}$.
Finally, the extension to arbitrary $n$ is straightforward: there are $n$ cycles, with $K_{k}=(i_{1}\to\ldots\to i_{k+1}\to i_{1})$ for all integer $31\leq k\leq n-$ 3, all of them generating the blocking pair ($i_{n},s_{1}$), and only $K_{1}$ introducing exactly this blocking pair. There are exactly three more cycles: $K_{n-2}=(i_{2}\to i_{3}\to\ldots\to i_{n-3}\to i_{2})$, $K_{n-1}=(i_{1}\to i_{6}\to i_{1})$, and $K_{n}=(i_{1}\to i_{n-1}\to i_{2}\to i_{1})$. Only one collection of cycles that leaves the graph acyclic is setwise justifiable envy minimal: $K_{1}$. However, the cycles $K_{n-2}$ and $K_{n-1}$ are compatible and lead to an allocation where $n-1$ students improve, and therefore any setwise justified envy minimal mechanism incurs in the largest improvement ratio possible, namely $\frac{n-1}{2}$. ∎
MIIDA also has important limitations beyond its lack of stability properties. The mechanism is manipulable and may have multiple solutions. Nonetheless, despite its limitations, we believe MIIDA serves as a useful mechanism to help us better understand the tradeoffs between efficiency and stability when improving over the celebrated DA mechanism.
@@ 6 Conclusion
We have shown that EADA and DA+TTC can be doubly dominated, improving as few as 2 students when $n-1$ could benefit, while generating more blocking pairs than larger improvements. MIIDA addresses this by maximizing the number of students who benefit, though our impossibility theorem proves this necessarily comes at the cost of losing EADA’s minimal justified envy guarantee.
An equally promising alternative would modify how EADA handles consent to minimize blocking pairs rather than maximize improvements. While such a mechanism might not improve as many students as MIIDA, it could nonetheless avoid double-dominance. This represents a complementary approach to navigating the efficiency-stability frontier that warrants future investigation.

### v3

@@ 4.2 Larger Improvements with Less Justified Envy
We find Corollary 1 useful because DA+TTC has been considered suboptimal regarding several axioms (Kesten, 2010; Troyan et al., 2020), yet it is not consistently doubly dominated by the more well-regarded EADA mechanism.
Second, the possibility of simultaneously achieving broader student coverage and fewer blocking pairs suggests that alternative efficient mechanisms that Pareto-dominate DA merit investigation. In particular, understanding the properties of such mechanisms—their computational complexity, laboratory performance, and incentive properties—represents an exciting area for future research. While EADA has deservedly received wide praise in the literature, our results point to unexplored opportunities within the class of mechanisms that Pareto-dominate DA that could consistently outperform EADA in this particular criterion of more beneficiaries with fewer blocking pairs.
Finally, we discuss the relationship between the impossibility result (Theorem 2) and the possibility result (Theorem 3). Both setwise and cardinal minimal instability capture legitimate concerns about mechanism design, though they reflect different policy goals. Setwise minimal instability focuses on the diversity of potential challenges to the allocation, recognizing that each blocking pair represents a distinct grievance that could motivate legal or political opposition. Cardinal minimal instability, by contrast, adopts a more utilitarian perspective that treats each blocking pair as imposing similar costs regardless of the specific students involved. While our impossibility result applies specifically to setwise minimally unstable mechanisms, this does not invalidate the setwise criterion but rather illuminates the fundamental trade-offs it entails. How much this limitation matters in practice depends critically on policymaker objectives: those prioritizing broad political consensus and legal robustness may find the setwise criterion compelling despite its constraints on student coverage, while those focused on maximizing total improvements or minimizing aggregate unfairness may prefer cardinal comparisons.
@@ 5 Conclusion
Our analysis establishes that setwise minimally unstable mechanisms, including EADA, face fundamental constraints on student coverage. We prove these mechanisms achieve the worst possible improvement ratio, potentially helping only $2$ students when n1n-1 could benefit. However, this limitation disappears under cardinal comparisons of blocking pairs: alternative mechanisms can simultaneously improve more students than EADA while generating fewer blocking pairs.
Our findings reveal that the choice between setwise and cardinal minimal instability represents an important design decision. While setwise minimally unstable mechanisms offer protection against diverse legal challenges, it does so at the cost of significantly reduced student participation in efficiency gains. The relatively underexplored class of mechanisms that Pareto-dominate DA contains promising alternatives for policymakers willing to prioritize broader student coverage over setwise stability constraints.